Find the standardized test statistic t for a sample with n = 12, x = 14.1, s = 1.3, and α = 0.05 if H0: μ ≥ 15. Round your answer to three decimal places.
Answer #1

Solution:

This is a left tailed test,

The test statistics,

t =( \bar x - \mu )/ (s /\sqrt{}n)

= ( 14.1 - 15 ) / ( 1.3 / \sqrt{} 12 )

= -2.398

P-value = 0.0177

The p-value is p = 0.0177 > 0.05, it is concluded that the null hypothesis is fail to rejected.